jabez177 (jabez177):

Check? :) Dayton is due north of Richmond, and Easton is 16 mi due east of Richmond. Dayton is 32 mi from Easton. How far apart are Dayton and Richmond? Round to the nearest tenth of a mile.

jabez177 (jabez177):

Let me draw a figure.

jabez177 (jabez177):



OpenStudy (aaronandyson):

Apply the Pythagorus Theorem.

OpenStudy (aaronandyson):

Let the desired distance by x Then x^2 + (16)^2 = (32)^2

jabez177 (jabez177):

16 x 16 = 256 32 x 32 = 1,024 1,024 - 256 = 768

jabez177 (jabez177):

square root of 768 = 27.71

jabez177 (jabez177):

so tenth of a mile will be 27.7 the answer is 27.7

jabez177 (jabez177):

right?

OpenStudy (aaronandyson):

Perfect!

OpenStudy (aaronandyson):

You already knew most of the problem and the best part is you did it yourself :)

jabez177 (jabez177):

I did put, "check" in the beginning. :P